OpenStudy (anonymous):

Estimate the area of a living room that measures 10.6 ft X 9.28 ft. _____

OpenStudy (anonymous):

PLEASE SHOW ME HOW TO SOLVE IT

OpenStudy (anonymous):

what are your options

OpenStudy (anonymous):

Estimate: 11*9 = 99 square feet Actual: 10.6*9.28 = 98.368 square feet

OpenStudy (anonymous):

90.0?

OpenStudy (anonymous):

what are the options

OpenStudy (anonymous):

there are no options.

OpenStudy (anonymous):

oh ok then

OpenStudy (anonymous):

how can i solve this?

OpenStudy (anonymous):

I gave you something didn't I Estimate: 11*9 = 99 square feet Actual: 10.6*9.28 = 98.368 square feet

OpenStudy (anonymous):

deoesn't this help

OpenStudy (anonymous):

yeah thx bye